Let x = red blood cell (RBC) count in millions per cubic millimeter of whole blood. Suppose that for healthy females, x has an approximately normal distribution with mean (weird looking u symbol) = 4.1 and standard deviation (weird looking o symbol) = 0.7. Convert the following z interval to an x interval.z < 2.9
Answer
a. x>6.13
b. x>-6.13
c. x<-6.13
d. x>-2.07
e. x<6.13
Could someone please show me how to solve this problem?

Answer by stanbon(75887)

Let x = red blood cell (RBC) count in millions per cubic millimeter of whole blood. Suppose that for healthy females, x has an approximately normal distribution with mean (weird looking u symbol) = 4.1 and standard deviation (weird looking o symbol) = 0.7. Convert the following z interval to an x interval.z < 2.9
------
You are given u = 4.1 and s = 0.7 and z = 2.9
---
The formula that relates x and z is x = z*s + u
Substitute and solve for "x":
x = 2.9*0.7 + 4.1 = 6.13
---
Ansswer: x > 6.13
========================
Cheers,
Stan H.
========================
Answer
a. x>6.13
b. x>-6.13
c. x<-6.13
d. x>-2.07
e. x<6.13